# Counterflow of lattice polarons in harmonically confined optical lattices

### v1

@@ Abstract
We study a mobile impurity in a one-dimensional harmonically confined optical lattice interacting repulsively with a bosonic bath. The behavior of the impurity across baths with superfluid and Mott-insulator domains is examined, including its full back-action effect on the bath. We characterize the bath-impurity phase diagram and reveal the appearance of a correlated counterflow phase, which we support with an analytical model for a mobile impurity-hole pair. This phase shows a combined insulator domain of unity filling but no independent domain of constant density. The transition to this phase features a sudden orthogonality catastrophe and the change of the shape of the impurity’s profile to that of a free particle in an infinite square well. The findings of this work suggest the appearance of unconventional counterflow in trapped imbalanced atomic mixtures.
@@
A rich platform for studying ultracold atomic impurities is optical lattices [22, 23, 24]. A decade ago, spin impurities interacting with a bosonic bath in a lattice chain were realized experimentally [25]. Such development motivated many theoretical studies of Bose lattice polarons, specially in the past few years [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. The latter include the study of polarons across the superfluid (SF) to Mott-insulator (MI) phase transition [45, 46, 47], where quantum critical effects become important. In this direction, it has been recently shown that the polaron properties are very sensitive to such transition [37, 43].
@@
In this letter, we consider a mobile impurity interacting repulsively with a bath of bosons, all trapped in a one-dimensional optical lattice. As depicted in Fig. 1(a), we add an external harmonic potential to study the effect of such a confinement. We examine the system across baths with SF and MI domains and reveal the appearance of a phase with counterflow order. Supercounterflow phases have been recently observed experimentally in binary MIs [61] after being predicted for many years [62, 63, 64, 65, 66, 67, 68, 69, 70]. In contrast with traditional supercounterflows [66], we find counterflow domains that do not show independent profiles with constant filling. We characterize the properties of this unconventional counterflow and present an analytical model that captures the main physics solely based on the formation of impurity-hole pairs [see Fig. 1(b)]. With this model, and by noting that the impurity behaves as a free particle in an infinite square well, we can predict the long-distance behavior of the anti-pair correlator.
@@
The properties of Bose lattice polarons are dictated by $t/U_{bb}$ and $U_{bI}/U_{bb}$ [37, 33]. However, the harmonic trap adds an additional scale, $\xi=d\sqrt{t/V_{\text{ho}}}$, where $d$ is the lattice spacing, which defines the bath’s characteristic density $\tilde{\rho}_{b}=dN_{b}/\xi$. The properties of the harmonically confined Bose-Hubbard model become invariant for constant $\tilde{\rho}_{b}$ [59]. Thus, this work examines the system for a varying range of $t/U_{bb}$ and $U_{bI}/U_{bb}$ but for a fixed density of $\tilde{\rho}_{b}=3.58$. This choice results in a bath that is completely SF for $t/U_{bb}\gtrsim 0.155$ and that has a central MI domain of unity filling for all $t/U_{bb}\lesssim 0.155$ [59]. We perform numerical simulations using the density matrix renormalization group (DMRG) method [71, 72, 73, 74]. See the Supplemental Material for more details.
Phase diagram and density profiles.– To examine the behavior of the system under different interaction strengths we compute the density profiles $n_{\sigma}(i)=\langle\hat{n}_{i,\sigma}\rangle$. We study the profiles as a function of the rescaled length $x_{i}=i\,d/\xi$ [59]. To characterize the phase diagram, we calculate the average position $\langle x_{\sigma}\rangle$ of each species and the size of the impurity’s cloud $\langle x^{2}_{I}\rangle$-$\langle x_{I}\rangle^{2}$, where
@@
We show average positions and cloud’s sizes in Fig. 2(a–c), which show well-defined phases. Profiles for a representative set of interactions are shown in panels (d–k).
@@
We first examine completely SF baths [$t/U_{bb}\gtrsim 0.155$]. When $U_{bI}<U_{bb}$ [region i in panels (a–c)], we observe that $\langle x_{\sigma}\rangle=0$ ($\sigma=b,I$), and thus both the bath and impurity are localized at the center of the trap. This is illustrated by the profiles (d) and (e), where the bath develops a small decrease in its density around $x_{i}=0$ due to the impurity’s repulsion. This decrease is only possible due to the compressibility of the SF bath. The extension of the impurity also shows a very small increase with $U_{bI}$, albeit difficult to observe in panel (c). Then, for $U_{bI}=U_{bb}$, the system simply behaves as an SF gas with $N_{b}+1$ bosons. This is illustrated by panels (f), where the impurity develops an enlarged profile, as captured by the large values in panel (c) for $U_{bI}=U_{bb}$. Finally, when $U_{bI}>U_{bb}$ [region ii], the impurity and bath develop finite average positions with opposite signs and a very large $\langle x_{I}\rangle$, signaling a phase separation between the species. This immiscibility is illustrated by panels (g), where the impurity is repelled to the right border of the trap  [^1] Note that the impurity localizes at only one border of the trap instead of developing a superposition on both sides. This is a feature of the DMRG calculations. In contrast, exact diagonalization calculations [81] do recover a symmetric solution with peaks at both borders.. The bath also moves slightly to the left, as illustrated by panel (b) where $\langle x_{b}\rangle\approx-0.05$. Note that in all cases the combined profiles $n_{b}+n_{I}$ [dashed purple lines in panels (d–g).2] show a Gaussian-like profile. Overall, the system’s behavior is similar to what is observed in non-lattice harmonically confined Bose polarons [51, 54].
@@
The situation changes when considering baths with an MI domain [$t/U_{bb}\lesssim 0.155$]. For small $U_{bI}$ [region iii], once again both the bath and impurity are centered in the trap ($\langle x_{\sigma}\rangle=0$). However, the profiles remain unchanged with $U_{bI}$. Indeed, in panel (h.1) the non-interacting profile (dotted curve) lies perfectly underneath the interacting one (solid curve). In turn, the bath [panel (h.2)] shows its MI domain of unity filling at the center, remaining undisturbed by the impurity. This behavior is due to the incompressibility of the MI domain. Turning to the case where $U_{bI}=U_{bb}$ [panels (j)], the system again behaves as a one-component one with $N_{b}+1$ bosons, thus developing an MI domain at the center, while the profile of the impurity is that of one particle in such a system. Then, for large $U_{bI}>U_{bb}$ [region iv], the system again undergoes a phase separation (large $\langle x_{I}\rangle$), as illustrated by panels (k). Due to the incompressibility, the immiscible MI bath centers at $\langle x_{b}\rangle\approx-0.09$, thus moving further away than an SF bath. Finally, an additional unexpected phase appears for intermediate interactions $U_{bI}$ (region v), which is absent in lattices with no harmonic confinement. This corresponds to a counterflow phase, which we examine in the following.
@@
Counterflow phase.– Region v of Fig. 2(a–c) develops notorious changes in the profiles, which are illustrated by Fig. 2(i). In panel (i.1), the interacting impurity’s profile (solid line) deviates strongly from the non-interacting one (dashed line), developing an enlarged profile. The latter is captured by the larger cloud’s size shown in Fig. 2(c). Both species also develop a finite but small $\langle x_{\sigma}\rangle$ [panels (a) and (b)], difficult to observe in the profiles, which indicates a small delocalization instead of a phase separation. Importantly, the bath decreases its density around the trap’s center instead of showing a domain of constant unity filling. Therefore, the stronger bath-impurity repulsion destroys the MI domain, making the bath compressible. For clarity, the profiles can be appreciated in greater detail in Fig. 3(a,b).
@@
A striking feature of this phase is that the sum of the profiles [dashed purple lines in Figs. 2(i.2) and 3(b)] form a domain of unity filling, signaling an insulating state. Thus, the bath and impurity form a combined insulator phase, even though, independently, each species does not show features of an MI domain. Indeed, the independent profiles do not show any plateau of constant density, only the sum of the two does. This behavior can be explained by the appearance of a counterflow state.
@@
Figure 3: Impurity’s (a), bath’s [solid red line in (b)] and combined [dashed purple line in (b)] profiles, and anti-pair correlator (c) as a function of $x_{i}$ for $t/U_{bb}=0.1$. Panels (a) and (b) consider $U_{bI}/U_{bb}=0.6$, while panel (c) considers the values indicated in the legend. The long-dashed black lines in (a–c) are a trigonometric fit for $U_{bI}/U_{bb}=0.6$ [see text] and the dash-dotted line in (c) is an exponential fit for $U_{bI}/U_{bb}=0.2$.
@@
Counterflow order is characterized by a slow algebraic decay of the anti-pair correlator [66]
$\mathcal{C}_{\text{SCF}}(i)=\langle\hat{a}_{0,b}\hat{a}^{\dagger}_{0,I}\hat{a}^{\dagger}_{i,b}\hat{a}_{i,I}\rangle.$ (3)
@@
$\mathcal{C}_{\text{SCF}}$ decays exponentially in all regions i to iv of the phase diagram. In Fig. 3(c), this rapid decay is illustrated by the results for $U_{bI}/U_{bb}=0.2$ and $U_{bI}/U_{bb}=0.4$ [region iii]. In contrast, for $U_{bI}/U_{bb}=0.6$ and $U_{bI}/U_{bb}=0.8$ [region v], $\mathcal{C}_{\text{SCF}}$ shows a noticeable slower decay. Moreover, the correlator is approximately constant for $x_{i}\lesssim 0.5$. We have found that this slow decay appears suddenly in region v, and thus, we can conclude that such a region corresponds to a counterflow phase.
In a counterflow state, each species tunnels in opposite directions [see Fig. 1(b)]. Importantly, counterflow phases are insulators, even though the particles can still move [61]. This correlated insulator state explains the appearance of the combined profile of unity filling $n_{b}+n_{I}=1$. However, while traditional binary counterflows also show individual profiles with plateaus of constant density [66], in our system the bath and impurity do not show such plateaus, as discussed. The latter makes the picture of an insulator particularly counterintuitive, as the impurity prefers to flow at the center of the lattice while the bath shows density peaks at $|x_{i}|\approx 1$. Nevertheless, the anti-pair correlator confirms the appearance of a counterflow.
We can now comment again on the sudden change of the impurity’s profile. In the non-counterflow phases, $n_{I}$ are Gaussian-like functions, where their widths are dictated by the harmonic trap. In contrast, the counterflow impurity develops an enlarged profile. Interestingly, we have found that the counterflow impurity’s profile can be fitted almost perfectly by the function
@@
where $n_{I}^{(0)}$ is the maximum value of the profile and $\ell_{\text{CF}}=2/(\xi n_{I}^{(0)})$ so that $\sum_{i}n_{I}^{(\text{CF})}=1$. Function (4) is shown in Fig. 3(a) by the long-dashed line. Additionally, we show $1-n_{I}^{(\text{CF})}$ in Fig. 3(b), which fits $n_{b}$ almost perfectly for $|x_{i}|\lesssim 1$. We have found a value of $\ell_{\text{CF}}\approx 3$ for $t/U_{bb}=0.1$ and $U_{bI}/U_{bb}=0.6$. Nevertheless, $\ell_{\text{CF}}$ remains almost constant in the whole counterflow phase, as the profiles show only small variations.
Naturally, Eq. (4) describes the ground-state density probability of a free particle in an infinite square well of width $\ell_{\text{CF}}$. This can be understood as an effective behavior. Indeed, the appearance of this length scale $\ell_{\text{CF}}$ stems from the strong correlations between the impurity and the bath, which dominate over the effect of the harmonic trap. Nonetheless, the appearance of this box-like profile is a striking new feature of the counterflow impurity, as one would instead expect a Gaussian profile.
@@
where $\sum_{i}\alpha^{2}_{i}=1$. With this wavefunction, the impurity’s profile reads $n_{I}(i)=\alpha^{2}_{i}$. Importantly, after some algebra, one obtains that the bath behaves as $n_{b}(i)=1-\alpha^{2}_{i}$, recovering the result $n_{b}+n_{I}=1$. In addition, the anti-pair correlator takes the form $\mathcal{C}_{\text{SCF}}(i)=\alpha_{0}^{*}\alpha_{i}$. Inspired by Eq. (4), from imposing $n_{I}(i)=n_{I}^{(0)}\cos^{2}(i\pi/M_{\text{CF}})$, one obtains a trigonometric dependence of the correlator
$\mathcal{C}_{\text{SCF}}(i)=n_{I}^{(0)}\cos(i\pi/M_{\text{CF}}),$ (7)
which predicts a slow decay. The long-dashed line in Fig. 3(c) shows function (7) with the constants taken from Eq. (4). Note that we have used $\ell_{\text{CF}}=dM_{\text{CF}}/\xi$ and $x_{i}\to(x_{i}-\langle x_{I}\rangle)$. The cosine gives a near-perfect agreement, and thus we can conclude that this simple impurity-hole model nicely predicts the behavior of $\mathcal{C}_{\text{SCF}}$.
Finally, we again stress that a counterflow is characterized by tunneling in opposite directions of the bath and impurity. While immiscible configurations also form impurity-hole pairs [42], they do not feature this correlated counter-tunneling.
@@
which measures the prevalence of the free impurity state in the interacting one, and the polaron energy
$E_{p}(U_{bI})=E(U_{bI})-E(U_{bI}=0).$ (9)
Here $E$ are ground-state energies, and thus $E_{p}$ measures the energy required to add the impurity to the bath. We show both quantities in Fig. 4.
@@
Firstly, the residue in panel (a) strongly marks the transition between the miscible regions i and ii with the other phases. A vanishing residue for $U_{bI}>U_{bb}$ is expected in immiscible configurations [51]. However, the residue also vanishes in the counterflow phase [region iii]. This reveals a orthogonality cathastrophe when counterflow order appears, indicating the breaking of the traditional polaron picture. Interestingly, in SF baths [region i] the residue decreases smoothly with increasing $U_{bI}$ [see circle markers in (c)], similar to other scenarios where polarons show orthogonality. In contrast, in baths with MI domains the residue remains constant and equal to one for small $U_{bI}$ [region ii], which we attribute to the incompressibility of the MI bath. Then, the residue vanishes abruptly at the transition to the counterflow phase [region v]. Such behavior is well-lustrated by the square markers in panel (c), where $Z$ shows a strong discontinuity. This sudden orthogonality is another feature of the counterflow phase, which could be probed experimentally via radio-frequency spectroscopy [77].
@@
Finally, the polaron energy [panel (b)] also shows distinct behaviors at the different phases, albeit with smooth transitions and no discontinuities. In all cases, $E_{p}$ increases with $U_{bI}$, to then saturate for large $U_{bI}$ [see panel (d)], agreeing qualitatively with related studies [29, 33]. Importantly, in baths with MI domains [see square markers in (d)] the polaron energy initially increases as $E_{p}=U_{bI}$ because the impurity only interacts with one boson per site. Then, in the counterflow phase, the polaron energy deviates from such dependency, showing a slower increase with $U_{bI}$ due to the change in the kinetic and interacting energies. These dependencies could be measured experimentally [78], providing an additional option to probe the counterflow phase.
@@
Conclusions.– We have shown that a mobile impurity interacting repulsively with a bosonic bath in a harmonically confined optical lattice shows a very rich behavior, including the appearance of an unconventional counterflow phase. This counterflow is characterized by the appearance of a combined domain of unity filling with long-range anti-pair correlations. Contrary to intuition, we found that in this phase the impurity behaves as a free particle moving in an effective infinite square well. The width of this well is much larger than the harmonic oscillator width associated with a free impurity, indicating that the effect of the bath-impurity correlations dominates over the harmonic confinement. The counterflow phase transition also features a sudden orthogonality catastrophe of the polaron quasiparticle. These features could be measured in current experiments with highly-imbalanced ultracold atomic mixtures, and also be used to probe the SF-MI transition in harmonically confined optical lattices. Overall, this work provides highly accurate results for the growing literature on lattice polarons. More importantly, the unconventional findings on counterflows could be further explored theoretically with imbalanced mixtures [79] and bipolarons [36, 39], and also experimentally following the recent realization of supercounterflows [61].
Acknowledgements.
@@
We thank J. Martorell for the useful discussions. F.I. acknowledges funding from ANID through FONDECYT Postdoctorado No. 3230023. This work has been funded by Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu). A.R.-F. acknowledges funding from MICIU through grant FPU20/06174. F.I. and L.M.-M. acknowledge additional funding from Instituto de Física UC. DMRG calculations were performed using the ITensor library [80].

### v2

@@ Abstract
We study a mobile impurity in a one-dimensional harmonically confined optical lattice interacting repulsively with a bosonic bath. The behavior of the impurity across baths with superfluid and Mott-insulator domains is examined, including its full back-action effect on the bath. We characterize the bath-impurity phase diagram and reveal the appearance of a correlated counterflow phase, which we support with an analytical model for a mobile impurity-hole pair. This phase shows an extended combined insulator domain of unity filling but no independent domain of constant density. The transition to this phase features a sudden orthogonality and the change of the shape of the impurity’s profile to that of a free particle in an infinite square well. The findings of this work suggest the appearance of unconventional counterflow in trapped imbalanced atomic mixtures.
@@
A rich platform for studying ultracold atomic impurities is optical lattices [22, 23, 24]. A decade ago, spin impurities interacting with a bosonic bath in a lattice chain were realized experimentally [25], motivating many theoretical studies of Bose lattice polarons in recent years [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. These include studies of polarons across the superfluid (SF) to Mott-insulator (MI) transition [45, 46, 47], where quantum critical effects become important. In this direction, it has been shown that the polaron properties are very sensitive to such transition [37, 43].
@@
In this letter, we consider a mobile impurity interacting repulsively with a bath of bosons, all trapped in a one-dimensional optical lattice. As depicted in Fig. 1(a), we add an external harmonic potential to study the effect of such a confinement. We examine the system across baths with SF and MI domains and reveal the appearance of a counterflow phase with anti-pair order. Supercounterflows have been recently realized experimentally in binary MIs [61] after being predicted for many years [62, 63, 64, 65, 66, 67, 68, 69, 70]. Thus, this work shows that counterflows can form for high atomic imbalances. Importantly, the impurity and bath become correlated over large distances, showing that it is not a localized effect. We characterize the properties of the counterflow state and present an analytical model that captures the main physics solely based on the formation of impurity-hole pairs [see Fig. 1(b)]. With this model, and by noting that the impurity behaves as a free particle in an infinite square well, we can explain the long-distance behavior of the anti-pair correlator.
@@
The properties of Bose lattice polarons are dictated by $t/U_{bb}$ and $U_{bI}/U_{bb}$ [37, 33]. However, the harmonic trap adds an additional scale, $\xi=d\sqrt{t/V_{\text{ho}}}$, where $d$ is the lattice spacing, which defines the bath’s characteristic density $\tilde{\rho}_{b}=dN_{b}/\xi$. The properties of the harmonically confined Bose-Hubbard model become invariant for constant $\tilde{\rho}_{b}$ [59]. Thus, this work examines the system for a varying range of $t/U_{bb}$ and $U_{bI}/U_{bb}$ but for a fixed density of $\tilde{\rho}_{b}=3.58$. This choice results in a bath that is completely SF for $t/U_{bb}\gtrsim 0.155$ and that has a central MI domain of unity filling for all $t/U_{bb}\lesssim 0.155$ [59]. We perform numerical simulations using the density matrix renormalization group (DMRG) method [71, 72, 73, 74] for $N_{b}=40$ and $M=60$. Nevertheless, our results are valid for any large $N_{b}$ due to the system’s invariance. See the supplemental material for additional details.
Phase diagram and density profiles.– To examine the behavior of the system under different interaction strengths, we compute the density profiles $n_{\sigma}(i)=\langle\hat{n}_{i,\sigma}\rangle$. We study the profiles as a function of the rescaled length $x_{i}=i\,d/\xi$ [59]. To characterize the phase diagram, we calculate the average position $\langle x_{\sigma}\rangle$ of each species and the size of the impurity’s cloud $\langle x^{2}_{I}\rangle$-$\langle x_{I}\rangle^{2}$, where
@@
We show average positions and clouds’ sizes in Fig. 2(a–c), which show well-defined phases. Profiles for a representative set of interactions are shown in panels (d–k).
@@
We first examine completely SF baths [$t/U_{bb}\gtrsim 0.155$]. When $U_{bI}<U_{bb}$ [region i in panels (a–c)], we observe that $\langle x_{\sigma}\rangle=0$ ($\sigma=b,I$), and thus both the bath and impurity are centered in the trap. This is illustrated by the profiles (d) and (e), where the bath develops a small decrease in its density around $x_{i}=0$ due to the impurity’s repulsion. This decrease is only possible due to the compressibility of the SF bath. The extension of the impurity also shows a very small increase with $U_{bI}$, albeit difficult to observe in panel (c). Then, for $U_{bI}=U_{bb}$, the system behaves as an SF gas with $N_{b}+1$ bosons. This is illustrated by panels (f), where the impurity develops an enlarged profile, as captured by the large values in panel (c) for $U_{bI}=U_{bb}$. Finally, when $U_{bI}>U_{bb}$ [region ii], the impurity and bath develop finite average positions with opposite signs and a very large $\langle x_{I}\rangle$, signaling a phase separation between the species. This immiscibility is illustrated by panels (g), where the impurity is repelled to the right border of the trap  [^1] Note that the impurity localizes at only one border of the trap instead of developing a superposition on both sides. This is a feature of the DMRG calculations. In contrast, exact diagonalization calculations [82] do recover a symmetric solution with peaks at both borders.. The bath also moves slightly to the left, as illustrated by panel (b) where $\langle x_{b}\rangle\approx-0.05$. Note that in all cases the combined profiles $n_{b}+n_{I}$ [dashed purple lines in panels (d–g).2] show a Gaussian-like profile. Overall, the system’s behavior is similar to what is observed in non-lattice harmonically confined Bose polarons [51, 54].
@@
The situation changes when considering baths with an MI domain [$t/U_{bb}\lesssim 0.155$]. For small $U_{bI}$ [region iii], once again both the bath and impurity are centered in the trap ($\langle x_{\sigma}\rangle=0$). However, the profiles remain unchanged with $U_{bI}$. Indeed, in panel (h.1), the non-interacting profile (dotted curve) overlaps with the interacting one (solid curve). In turn, the bath [panel (h.2)] shows its MI domain at the center, remaining undisturbed by the impurity. This behavior is due to the incompressibility of the MI domain. Turning to the case where $U_{bI}=U_{bb}$ [panels (j)], the system again behaves as a one-component one with $N_{b}+1$ bosons, and so the impurity’s profile is that of one particle in such a system. Then, for large $U_{bI}>U_{bb}$ [region iv], the system undergoes a phase separation (large $\langle x_{I}\rangle$), as illustrated by panels (k). Due to its incompressibility, the MI bath centers at $\langle x_{b}\rangle\approx-0.09$, thus moving further away than an SF bath. Finally, an additional phase with counterflow order appears for intermediate interactions (region v), which we examine next.
@@
Counterflow phase.– Region v of Fig. 2(a–c) develops notorious changes in the profiles, which are illustrated by Fig. 2(i). In panel (i.1), the interacting impurity’s profile (solid line) deviates strongly from the non-interacting one (dashed line), developing an enlarged profile. The latter is captured by the larger cloud’s size shown in Fig. 2(c). Both species also develop a finite but small $\langle x_{\sigma}\rangle$ [panels (a) and (b)], difficult to observe in the profiles. Importantly, the bath decreases its density around the trap’s center instead of showing a domain of constant unity filling. Therefore, the stronger bath-impurity repulsion destroys the MI domain, making the bath compressible. For clarity, the profiles can be appreciated in greater detail in Fig. 3(a,b).
@@
A striking feature of this phase is that the sum of the profiles [dashed purple lines in Figs. 2(i.2) and 3(b)] form a domain of unity filling, signaling an insulating state. Thus, the bath and impurity form a combined insulator phase, even though independently, each species does not show features of an MI domain. Indeed, the independent profiles do not show any plateau of constant density, only the sum of the two does. This behavior can be explained by the appearance of a counterflow state.
@@
Figure 3: Impurity’s (a), bath’s [solid red line in (b)] and combined [dashed purple line in (b)] profiles, and anti-pair correlator (c) as a function of $x_{i}$ for $t/U_{bb}=0.1$. Panels (a) and (b) consider $U_{bI}/U_{bb}=0.6$, while panel (c) considers the values indicated in the legend. The long-dashed black lines in (a–c) are functions (4) and (7), and the dash-dotted line in (c) is an exponential fit for $U_{bI}/U_{bb}=0.2$.
@@
Long-range counterflow order is characterized by a slow algebraic decay of the anti-pair correlator [66]
$\mathcal{C}_{\text{AP}}(i)=\langle\hat{a}_{0,b}\hat{a}^{\dagger}_{0,I}\hat{a}^{\dagger}_{i,b}\hat{a}_{i,I}\rangle.$ (3)
@@
$\mathcal{C}_{\text{AP}}$ decays exponentially in all regions i to iv of the phase diagram. In Fig. 3(c), this rapid decay is illustrated by the results for $U_{bI}/U_{bb}=0.2$ and $U_{bI}/U_{bb}=0.4$ [region iii]. In contrast, for $U_{bI}/U_{bb}=0.6$ and $U_{bI}/U_{bb}=0.8$ [region v], $\mathcal{C}_{\text{AP}}$ shows a noticeable slower decay. Moreover, the correlator is approximately constant for $x_{i}\lesssim 0.5$. We have found that this slow decay appears suddenly in region v, and thus, we can conclude that such a region corresponds to a counterflow phase. These results confirm that counterflow order can appear in highly imbalanced configurations, thus not being limited to balanced or near-balanced mixtures.
In a counterflow state, each species tunnels in opposite directions [see Fig. 1(b)]. Importantly, counterflows are insulators, even though the particles can still move [61]. This correlated insulator state explains the appearance of the combined profile of unity filling $n_{b}+n_{I}=1$. However, while balanced counterflows show individual profiles with plateaus of constant density [66], in our system, the bath and impurity do not show such plateaus, as discussed. This is unexpected, as one would naively predict plateaus with filling of $N_{b}/(N_{b}+1)$ and $1/(N_{b}+1)$. These unusual profiles make the picture of an insulator particularly counterintuitive, as the impurity prefers to flow at the center of the lattice while the bath shows density peaks at $|x_{i}|\approx 1$. Nevertheless, the behavior of $\mathcal{C}_{\text{AP}}$ confirms the appearance of a counterflow.
The slow decay of $\mathcal{C}_{\text{AP}}$ and the large extension of $n_{I}$ also demonstrate that the counterflow state forms over large distances. Therefore, this single impurity suddenly forms particle-hole pairs with the whole MI domain, indicating that it is not a localized effect.
The shape of the impurity’s profile is particularly interesting. In the non-counterflow phases, $n_{I}$ are Gaussian-like functions, where their widths are dictated by the harmonic trap. In contrast, we have found that the counterflow impurity’s profile takes the form
@@
where $n_{I}^{(0)}$ is the maximum value of the profile and $\ell_{\text{CF}}=2/(\xi n_{I}^{(0)})$ so that $\sum_{i}n_{I}^{(\text{CF})}=1$. Function (4) is shown in Fig. 3(a) by the long-dashed line. In turn, $1-n_{I}^{(\text{CF})}$ is shown in Fig. 3(b), which agrees with $n_{b}$ almost perfectly for $|x_{i}|\lesssim 1$. We have found a value of $\ell_{\text{CF}}\approx 3$ for $t/U_{bb}=0.1$ and $U_{bI}/U_{bb}=0.6$, even though $\ell_{\text{CF}}$ is roughly constant in the whole counterflow phase.
Eq. (4) describes the ground-state density probability of a free particle in an infinite square well of width $\ell_{\text{CF}}$. This can be understood as an effective behavior. The appearance of this length scale $\ell_{\text{CF}}$ stems from the strong correlations between the impurity and the bath, which dominate over the effect of the harmonic trap.
@@
where $\sum_{i}|\alpha_{i}|^{2}=1$. With this wavefunction, the impurity’s profile reads $n_{I}(i)=|\alpha_{i}|^{2}$. Importantly, after some algebra, one obtains that the bath behaves as $n_{b}(i)=1-|\alpha_{i}|^{2}$, recovering the result $n_{b}+n_{I}=1$. In addition, the anti-pair correlator takes the form $\mathcal{C}_{\text{AP}}(i)=\alpha_{0}^{*}\alpha_{i}$. By using Eq. (4), one obtains a trigonometric dependence
$\mathcal{C}_{\text{AP}}(i)=n_{I}^{(0)}\cos(i\pi/M_{\text{CF}}),$ (7)
which predicts a slow decay of $\mathcal{C}_{\text{AP}}$. The long-dashed line in Fig. 3(c) shows function (7) with the constants taken from Eq. (4). Note that we have used $\ell_{\text{CF}}=dM_{\text{CF}}/\xi$ and $x_{i}\to(x_{i}-\langle x_{I}\rangle)$. The cosine gives a near-perfect agreement, and thus we can conclude that this simple impurity-hole model nicely explains the behavior of $\mathcal{C}_{\text{AP}}$. Here we mention that while a Gaussian fit can qualitatively describe the profile of the impurity, its exponential nature cannot describe the slow decay of $\mathcal{C}_{\text{AP}}$.
Finally, the MI-counterflow transition studied here does not appear in lattices with no harmonic trap. In such lattices, an MI phase is only possible for integer filling $N_{b}=\nu M$, with $\nu$ an integer. However, a counterflow state requires that $N_{b}+N_{I}=\nu$. Thus, the harmonic trap is required to find the rich physics presented here.
@@
where $|\Psi\rangle$ is the ground-state wavefunction. The residue measures the prevalence of the free impurity state in the interacting one. We also calculate the polaron energy
$E_{p}(U_{bI})=E(U_{bI})-E(U_{bI}=0),$ (9)
where $E$ are ground-state energies, and thus $E_{p}$ measures the energy required to add the impurity to the bath. We show both quantities in Fig. 4.
@@
Firstly, the residue in panel (a) strongly marks the transition between the miscible regions (i and iii) and the other phases (ii, iv, and v). A vanishing residue for $U_{bI}>U_{bb}$ is expected in immiscible configurations [51]. However, the residue also vanishes in the counterflow phase. This orthogonality suggests an orthogonality catastrophe in the thermodynamic limit when counterflow order appears, indicating the breaking of the traditional polaron picture. We can understand this orthogonality by realizing that, due to the appearance of sites with holes, the Fock states that describe a counterflow configuration become orthogonal to those of an unperturbed MI bath.
@@
Interestingly, in SF baths [region i] the residue decreases smoothly with increasing $U_{bI}$ [see circle markers in (c)], similar to other scenarios where polarons show orthogonality. In contrast, in baths with MI domains, the residue remains constant and equal to one for small $U_{bI}$ [region ii], which we attribute to the incompressibility of the MI bath. Then, the residue vanishes rapidly at the transition to the counterflow phase [region v]. This behavior is well-illustrated by the square markers in panel (c). Here we stress that while the numerical calculations do not show a fully discontinuous residue, the discontinuity becomes more abrupt with a larger number of particles. This orthogonality is another feature of the counterflow phase, which could be probed experimentally via radio-frequency spectroscopy [77].
@@
Finally, the polaron energy [panel (b)] also shows distinct behaviors at the different phases, albeit with smooth transitions and no discontinuities. In all cases, $E_{p}$ increases with $U_{bI}$, to then saturate for large $U_{bI}$ [see panel (d)], agreeing qualitatively with related studies [29, 33]. Importantly, in baths with MI domains [see square markers in (d)], the polaron energy initially increases as $E_{p}=U_{bI}$ because the impurity only interacts with one boson per site. Then, in the counterflow phase, the polaron energy deviates from such dependency, showing a slower increase with $U_{bI}$ due to the change in the kinetic and interacting energies. These dependencies could be measured experimentally [78], providing an additional option to probe the counterflow phase.
@@
Conclusions.– We have shown that a mobile impurity interacting repulsively with a bosonic bath in a harmonically confined optical lattice shows a very rich behavior, including the appearance of an unconventional counterflow phase. This counterflow is characterized by the appearance of a combined domain of unity filling with long-range anti-pair correlations. Contrary to intuition, we found that in this phase the impurity behaves as a free particle moving in an effective infinite square well. The counterflow phase transition also features a sudden orthogonality between the interacting and non-interacting ground states. These features could be measured in current experiments with highly-imbalanced ultracold atomic mixtures, and also be used to probe the SF-MI transition in harmonically confined optical lattices. The findings on counterflows could be further explored theoretically with imbalanced mixtures [79] and bipolarons [36, 39], and also experimentally following the recent realization of supercounterflows [61]. Additionally, the revealed counterflow state of impurities could be used to probe and simulate broader anti-pair behaviors of polarons, such as in analog models [80].
@@
Acknowledgments.– We thank J. Martorell for the useful discussions. F.I. acknowledges funding from ANID through FONDECYT Postdoctorado No. 3230023. This work has been funded by Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu). A.R.-F. acknowledges funding from MICIU through grant FPU20/06174. F.I. and L.M.-M. acknowledge additional funding from Instituto de Física UC. DMRG calculations were performed using the ITensor library [81].